OpenStudy (anonymous):

Solve for x: \frac{ 1 }{ 27 }=9^{7x+3}

OpenStudy (anonymous):

I will adjust the format.. one moment..

OpenStudy (anonymous):

make the bases of 2 sides equal to each other then u can equate the powers and find the value of x.

OpenStudy (anonymous):

\[1/27=9^7x+3\]

OpenStudy (anonymous):

u mean\[\frac{ 1 }{ 27 }=9^{7x+3}\]

OpenStudy (anonymous):

i get 3^3?

OpenStudy (anonymous):

yes, sorry

OpenStudy (anonymous):

np :) ... yeah u r close but note that\[\frac{1}{27}=\frac{1}{3^3}=3^{-3}\]

OpenStudy (anonymous):

oh of course! got that..

OpenStudy (anonymous):

and then make the 9 = 3^2?

OpenStudy (anonymous):

yeah...quite right

OpenStudy (anonymous):

\[3^{-3}=3^{14x+6}\]

OpenStudy (anonymous):

Right...

OpenStudy (anonymous):

now go for solving for x

OpenStudy (anonymous):

\[-\frac{ 9 }{ 14 }\]

OpenStudy (anonymous):

you rock! thanks for the help.. the -3 is what got me! :)

OpenStudy (anonymous):

yw :)